please help me i am just not getting it .
i have to fill out the blank with approriate inequality symbol i have to show work and support my conclusion.
3/4______13/16

simplify -5[1-7](-6)+1
Evaluate 4x-y when x= 3 and y=_7.

Answer by stanbon(48546)

3/4______13/16
Change 3/4 to ?/16 so you can compare the number to 13/16
(3/4)(4/4)=12/16
So your problem compares 12/16 to 13/16
so you need "<" between the numbers
3/4=12/16<13/16
--------------
simplify -5[1-7](-6)+1
=-5*-6(-6)
=30*-6
=-180
---------
Evaluate 4x-y when x= 3 and y=_7.
=4*3--7
=12+7
=19
-----------
By the way, the rule for evaluating expressions like this is:
1st: evaluate the parentheses individually.
2nd: Use the distributive rule to get rid of the parentheses
3rd: Evaluate from left to right taking care of * and / as they occur.
4th: Evaluate from left to right taking care of + and - as they occur.
---------
Ex : -3(7-6/3+5)/2/(-4-7/14)
= -3(7-2+5)/2/(-4-1/2)
= -3(10)/2/(-(9/2))
= -30/2/(-9/2)
= -15/(-9/2)
= 30/9
= 10/3
Cheers,
Stan H.